OpenStudy (anonymous):

Evaluate 2x2 − 5y when x = -3 and y = 4.

OpenStudy (anonymous):

Is that 2x^2? (the x is squared)?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-2 the answer

OpenStudy (anonymous):

Assuming it is \[2x ^{2}-5y\] You evaluate as \[(2 \times-3 ^{2}) - (5\times4)\] Which is the same as \[(2 \times-3\times-3) - (5\times4)\] Which is the same as \[(2 \times9) - (5\times4)\]

OpenStudy (anonymous):

Yup, you got it MzR, good job.

OpenStudy (anonymous):

Which, as has been posted previously, is 18-20=-2

OpenStudy (anonymous):

thanks both of u...:)

OpenStudy (anonymous):

u there